OpenStudy (anonymous):

real-world examples of exponential growth and decay Can someone help me with this???

OpenStudy (andriod09):

hehe. i just finished this unit, there are multiple ones in science with how things like radioactive elements like uranium.

OpenStudy (anonymous):

Can you give me an example of one please??

OpenStudy (andriod09):

I just did. the element uranium as it decays, it decays exponentially. like if there is a 10g piece, it will decay 2g every week, then it would be 8g, then 6g, then 4g, then 2g, then it gets smaller and smaller and smaller. too many numbers

OpenStudy (anonymous):

So would that be the answer did you solve it @andriod09

OpenStudy (anonymous):

??? Did you solve it???

OpenStudy (klimenkov):

The probability to die at moment \(t\) (seconds, minutes, hours, months ... from the birth) is \(p(t)=1-e^{-\alpha t}\) where \(\alpha\) is constant.